OpenStudy (anonymous):

Solve the equation for y: 4x + 3y = -12.

OpenStudy (anonymous):

4x+3y=-12/ :3 (4/3)x +y=-4 y=-(4/3)x-4

OpenStudy (anonymous):

y = -4 - 4/3 x? is that right?